OpenStudy (anonymous):

the difference between two numbers is 8 if the lesser number is m what is the greater number

OpenStudy (anonymous):

g-m=8 g=8+m

OpenStudy (anonymous):

let the greater no. be x then according to the question: x-m=8 therefore x=8+m

OpenStudy (skullpatrol):